OpenStudy (anonymous):

a^2 + b^2 = c^2 a=7 b=7 c=?

OpenStudy (cwrw238):

7^2 + 7^2 = 49 + 49 can you continue

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\]\[7^2 + 7^2 = c^2\]\[49 + 49 = c^2\]\[98 = c^2\]Find the square root of both sides.